OpenStudy (anonymous):

Perform the indicated operation. 3k/k-2 + 6/2-k = 1) 3 2) -3 3) 3k+6/k-2 4) 3k+6/k+2

OpenStudy (anonymous):

\[\frac{3k}{k-2}+\frac{6}{2-k}\] can be rewritten as \[\frac{3k}{k-2}-\frac{6}{x-2}\] now that the denominators are the same, subtract in the numerator

OpenStudy (anonymous):

once you get \[\frac{3k-6}{k-2}\] you can factor the numerator, then cancel \[\frac{3k-6}{k-2}=\frac{3(k-2)}{k-2}=\frac{3\cancel{(k-2)}}{\cancel{k-2}}=3\]